Solve the system by using the substitution method. x-2y= 10, 2x-4y= 16
Answer by oberobic(2304)

x-2y= 10
2x -4y = 16
.
multiply the first equation by 2
.
2x -4y = 20
2x -4y = 16
.
That's impossible.